Bonds of Zello Corporation with a par value of $1,000 sell for $1,080, mature in 18 years, and have a 7% annual coupon rate paid annually, what is the yield to maturity? what is the current yield? what is the capital gains (loss) yield?

please show step by step solution.for full credit, Not excel workings. Thank you!!!
Answer #1

Using financial calculator
Input: FV= 1000,

PV = -1080

N = 18

PMT = 7%*1000 = 70

Solve for I/Y as 6.25

Hence YTM = 6.25%

2: Current yield = Annual coupon/Price

= 70/1080 = 6.48%

3: Capital gain yield = (P1-P0)/P0

= (1000-1080)/1080

= -7.41%